OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, 8>, v = <-4, 8>

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Oh gees...there's a formula for that. I don't remember it. I think it involves the dot product which I also don't remember.

OpenStudy (mertsj):

\[\cos \theta=\frac{(-5)(-4)+(8)(8)}{\sqrt{(-5)^2+(8)^2)}\sqrt{(-4)^2+(8)^2}}\]

OpenStudy (turingtest):

The formula in general is\[\vec u\cdot\vec v=\|\vec v\|\|\vec u\|\cos\theta_{u,v}\]which when solved for \(\theta\) gives the same result as @Mertsj gave

OpenStudy (anonymous):

I got 89 degrees but it's incorrect :( @TuringTest @Mertsj

OpenStudy (mertsj):

Better check your arithmetic.